2. A histogram of monthly wages for part-time employees is shown below (densities are marked in parentheses) Nobody earned more than $1,000 a month. The block over the class interval from $300 to $600 is missing. What is the area and height of the bar? (write answers in the blanks; 3 points each) (14) Area 14 4乙 Height- 7 1 2 3 45 6 7 89 10 Wages (hundreds of dollars) Page 1 of 2
Answer #1

The heights of the blocks are given in the DENSITY SCALE of (% per $100). Hence, the areas of the blocks come out in %s.

The area of first block is = 1(wages) * 7(%per $`100) = 1*7 = 7%

The area of second block is = 2(wages) * 14(%per $`100) = 2*14 = 28%

The area of last block = 4(wages)*5(% per $100) = 4*5 =20%

The 3 blocks together gives = (7%+28%+20%) = 55%

Therefore the area of missing block is 100%-55% =45%

Height of missing block is = 45%/3(wages) = 15(% per $100)

Answer: Area = 45

Height = 15